The ratio of Brand "Y" golf balls sold to Brand "X" golf balls sold was 4 to 5. If there were 22 fewer Brand "Y" golf ablls than Brand "X" golf balls, then how many Brand "X" golf balls were sold?
Answer by stanbon(48568)

The ratio of Brand "Y" golf balls sold to Brand "X" golf balls sold was 4 to 5. If there were 22 fewer Brand "Y" golf ablls than Brand "X" golf balls, then how many Brand "X" golf balls were sold?
-------------
Let number of brand Y balls be 4x
Then number of brand x balls is 5x
---------
EQUATION:
# of y balls + 22 = # of x balls
4x+22 = 5x
x=22
--------
# of x balls = 5x = 5*22 = 110
============
Cheers,
Stan H.